# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v4

@@ 4 Theory
Theorem 1.
The quantity $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is SO(3)-invariant, where $\mathbf{Q}$ is the simplified representation (without superscripts) of $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
Theorem 2.
@@
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}},\quad\text{subject to}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents arbitrary differentiable non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
The proofs of the two theorems above are presented in Appendix B.
Figure 1: Methodological framework for learning SO(3)-equivariant representations with non-linear expressiveness to regress complex SO(3)-equivariant targets.
@@ 5 Method
As shown in Fig. 1, taking $z$ in Eq. 5 serving as the bridge between Theorem 1 and Theorem 2, we propose a general method for learning non-linear representations that satisfy the SO(3)-equivariance property outlined in Eq. 4. The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning $z$, i.e., the SO(3)-invariant internal representations of $g_{\text{nonlin}}(\cdot)$ in Theorem 2. Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities. At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations for inferring the SO(3)-equivariant regression target.
@@ 5.1 Encoding and Decoding Framework
Our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a deep encoding framework. For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant backbone encoder, e.g., the encoders from DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b) which is composed of recently developed equivariant operators (Thomas et al., 2018; Geiger & Smidt, 2022) like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics (Schrodinger, 1926), or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer. Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$, $s_{nonlin}(\cdot)$ are the non-linear functions defined in Eq. 5. The function $s_{nonlin}(\cdot)$ can be implemented as any differentiable non-linear neural network module, such as feed-forward layers with non-linear activation functions like $SiLU$ and normalization operations like $Layernorm$.
For an expressive representation of a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\{\textbf{f}^{(k)l_{1}},\textbf{f}^{(k)l_{2}},\textbf{f}^{(k)l_{3}},...\}$ at multiple degrees, i.e., $L^{(k)}=\{l_{1},l_{2},l_{3},...\}$, where some components share the same degree while others differ. In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq. 5, to accommodate various components of degrees. Moreover, in the context of neural networks, introducing learnable parameters $W$ and more feature channels $C$ may improve the model capacity. Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq. 5 can be expanded as $\text{CGDecomp}_{ext}(\cdot)$, and its outputs can be expanded as $\textbf{u}^{(k)}$, as shown in Eq. 6:
@@
$u_{c}^{(k)}$ $=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{l_{i},l_{j}\in L^{(k)},l_{i}=l_{j}}W^{c}_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$
@@
[…]
@@ 5.2 Training
$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=Error(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\,\,\,\,loss_{T}=Error(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$. In order to prevent the numerical disparity between the two loss terms and stabilize the training for both of the SO(3)-equivariant and SO(3)-invariant branches, we apply $\mu(loss_{Q},loss_{T})$, i.e., a coefficient to regularize the relative scale of the two loss terms:
$\mu(loss_{Q},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{Q}}{loss_{T}})$ (8)
where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale between the two loss terms and should not itself be a source of training gradients, otherwise it would counteract the gradients from $loss_{T}$ in Eq. 7. All of the encoding and decoding modules are trained jointly by Eq. 7.
For the implementation details of our method, including the specific design of the network modules, parameter settings, and training specifics, please refer to Appendix D.
@@ 6.1 Experimental Conditions
We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-stable (QS), and QH9-dynamic (QD). The first six databases, consisting of periodic crystalline systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023). The last two databases are from the QH9 benchmark series (Yu et al., 2023a), composed of molecular systems with elements like C, H, O, N and F. These databases present diverse and complex challenges to the generality of a regression model. Regarding $MG$, $MM$, and QD, as their samples are prepared from an temperature environment at three-hundred Kelvin, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction. For $BG$, $BB$, $BT$, and $BS$, the twisted structures, with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets. Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling. For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and testing sets, ensuring that the atom number of samples do not overlap across the three subsets. For the $QD$ database, the ’mol’ strategy provided by Yu et al. (2023a) is applied to split the training, validation, and testing sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets. The ’mol’ and ’ood’ strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories, respectively. Detailed statistic information of these databases can be found in the Appendix C. Implementation details of our method for experiments on these databases are presented in Appendix D.
@@
We use a comprehensive set of metrics to deeply evaluate the performance of deep learning electronic-structure Hamiltonian prediction models. The metrics can be broadly categorized into two major aspects:
@@
• Metrics on accuracy performance. On the databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023), we follow Yin et al. (2024) to adopt a set of Mean Absolute Error (MAE) metrics between predicted and ground truth Hamiltonians, including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$). These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices. On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper (Yu et al., 2023a), including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians. $\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
• Metrics on acceleration ratios for the convergence of traditional DFT algorithms. Despite the increasing ability of deep learning models to independently handle more electronic-structure computation tasks, there are still applications with extremely high numerical precision requirements and very low tolerance for error, where traditional DFT algorithms must perform the final calculations. In such cases, the predictions from deep models can be used as initial matrices to accelerate the convergence of traditional DFT algorithms. We evaluate the acceleration ratios brought by the proposed method for the convergence of classical DFT algorithms implemented by PySCF (Sun et al., 2018). Specifically, we adopt the two metrics defined in Yu et al. (2023a): the achieved ratio calculates the number of DFT optimization steps taken when initializing with the Hamiltonian matrices predicted by the deep model compared to using initial guess methods like minao and 1e. The error-level ratio measures the number of DFT optimization steps required, starting from random initialization, to reach the same error level as the deep model’s predictions, relative to the total number of steps in the DFT process.
Table 3: The acceleration ratios of QHNet and QHNet+TraceGrad for DFT calculation. Both models are evaluated on a set of $50$ molecules chosen by Yu et al. (2023a), with the mean and standard deviation of the metrics across these samples reported. $\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance.
Methods Training DFT Metric Ratio
databases initialization
QHNet (Baseline) QS 1e Achieved ratio $\downarrow$ $0.400\pm 0.030$
Error-level ratio $\uparrow$ $0.620\pm 0.037$
minao Achieved ratio $\downarrow$ $0.715\pm 0.033$
Error-level ratio $\uparrow$ $0.406\pm 0.021$
QD 1e Achieved ratio $\downarrow$ $0.512\pm 0.138$
Error-level ratio $\uparrow$ $0.622\pm 0.048$
minao Achieved ratio $\downarrow$ $0.882\pm 0.217$
Error-level ratio $\uparrow$ $0.406\pm 0.066$
Error-level ratio $\uparrow$ $\textbf{0.685}\pm 0.037$
minao Achieved ratio $\downarrow$ $\textbf{0.647}\pm 0.061$
Error-level ratio $\uparrow$ $\textbf{0.466}\pm 0.035$
QD 1e Achieved ratio $\downarrow$ $\textbf{0.440}\pm 0.101$
Error-level ratio $\uparrow$ $\textbf{0.645}\pm 0.046$
minao Achieved ratio $\downarrow$ $\textbf{0.761}\pm 0.167$
Error-level ratio $\uparrow$ $\textbf{0.435}\pm 0.052$

### v5

@@ 4 Theory
Theorem 4.1.
The quantity $\mathbf{T}=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ is SO(3)-invariant, where $\mathbf{H}$ is the simplified representation (without superscripts) of the basic block $\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}$ of Hamiltonian matrix defined in Appendix B, and $\dagger$ denotes the conjugate transpose operation, $tr(\cdot)$ is the trace operation.
Theorem 4.2.
@@
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}},\quad$ (1)
$\text{subject to}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the SO(3)-invariant component corresponding to degree 0, $s_{nonlin}(\cdot)$ represents arbitrary differentiable non-linear neural modules, $u$ and $z$ are SO(3)-invariant features, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
The proofs of the two theorems above are presented in Appendix D.
Remark 4.3.
In physics, dimensions describe the fundamental properties of physical quantities. The seven base dimensions are mass $[M]$, time $[T]$, length $[L]$, electric current $[I]$, temperature $[\Theta]$, amount of substance $[N]$, and luminous intensity $[J]$. Other physical quantities can be derived from these base dimensions. For instance, the dimension of energy $E$ is given by $\dim E=ML^{2}T^{-2}$. The quantities $\mathbf{H}$ and $\mathbf{T}$ presented in Theorem 4.1 correspond to the physical dimensions of $\dim E$ and $\dim E^{2}$, respectively, where their units can be taken as meV and (meV)2. These two physical dimensions are related through: $\dim E=\frac{\dim E^{2}}{\dim E}$. As we developed in Section 5, in the feature space, the SO(3)-invariant feature $z$ is used to regress $\mathbf{T}$, and the SO(3)-equivariant features $\mathbf{f}$ and $\mathbf{v}$ are used to regress $\mathbf{H}$. Our gradient-based mechanism $\mathbf{v}=\frac{\partial z}{\partial\mathbf{f}}$ in Theorem 4.2 naturally preserves the dimensional correspondence between $\mathbf{H}$ and $\mathbf{T}$. On the other hand, the existing gated activation function (Weiler et al., 2018), which can be expressed as $\mathbf{v}=z\cdot\mathbf{f}$, does not enforce such a dimensional correspondence and could not offer a strong guarantee of a balanced coupling between the SO(3)-invariant and SO(3)-equivariant representations. In contrast, our approach, with adherence to the dimensional correspondence between $\mathbf{H}$ and $\mathbf{T}$, ensures that the learned features are more physically grounded and expressive, effectively capturing the underlying relationships between SO(3)-invariant and SO(3)-equivariant quantities. See Appendix H for an empirical study comparing between the gradient-based mechanism and the gated mechanism.
Figure 1: The proposed framework for learning SO(3)-equivariant representations with strong non-linear expressiveness to regress the SO(3)-equivariant electronic-structure Hamiltonian.
@@ 5 Method
As shown in Fig. 1, taking $z$ in Eq. (1) serving as the bridge between Theorem 4.1 and Theorem 4.2, we propose a method for learning expressive non-linear representations that satisfy the SO(3)-equivariance property outlined in Eq. (10). The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ introduced in Theorem 4.1 as a supervisory signal for learning $z$, i.e., the SO(3)-invariant internal representations of $g_{\text{nonlin}}(\cdot)$ in Theorem 4.2. Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities. At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations for inferring the SO(3)-equivariant electronic-structure Hamiltonians.
@@ 5.1 Encoding and Decoding Framework
Our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a deep encoding framework. For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant backbone encoder, e.g., the encoders from DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b) which is composed of recently developed equivariant operators (Thomas et al., 2018; Geiger & Smidt, 2022) like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics (Schrodinger, 1926), or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer. Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$, $s_{nonlin}(\cdot)$ are the non-linear functions defined in Eq. (1). The function $s_{nonlin}(\cdot)$ can be implemented as any differentiable non-linear neural network module, such as feed-forward layers with non-linear activation functions like $SiLU$ and normalization operations like $Layernorm$.
For an expressive representation of a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\{\textbf{f}^{(k)l_{1}},\textbf{f}^{(k)l_{2}},\textbf{f}^{(k)l_{3}},...\}$ at multiple degrees, i.e., $L^{(k)}=\{l_{1},l_{2},l_{3},...\}$, where some components share the same degree while others differ. In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq. (1), to accommodate various components of degrees. Moreover, in the context of neural networks, introducing learnable parameters $W$ and more feature channels $C$ may improve the model capacity. Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq. (1) can be expanded as $\text{CGDecomp}_{ext}(\cdot)$, and its outputs can be expanded as $\textbf{u}^{(k)}$, as shown in Eq. (2):
@@
$u_{c}^{(k)}$ $=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)$
$=\sum_{l_{i},l_{j}\in L^{(k)},l_{i}=l_{j}}W^{c}_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$
@@
We follow previous literature (Gong et al., 2023; Yu et al., 2023b) to send the features from the last layer of the encoder, i.e., $\textbf{o}^{(K)}$ in our framework, into the SO(3)-equivariant decoder to regress the predictions of $\mathbf{H}$. For this, we can directly utilize the mature design of the SO(3)-equivariant decoders in DeepH-E3 or QHNet. The new part in the decoding phase introduced by our method is the SO(3)-invariant decoder, consisting of feed-forward layers taking $\textbf{z}=[\textbf{z}^{(1)},\ldots,\textbf{z}^{(k)},\ldots,\textbf{z}^{(K)}]$ as the input to predict $\mathbf{T}$.
@@ 5.2 Training
$\begin{split}\min_{\theta}loss&=loss_{H}+\mu(loss_{H},loss_{T})\cdot loss_{T},\\ loss_{H}&=Error(\widehat{\mathbf{H}},\mathbf{H}^{*}),\,\,\,\,loss_{T}=Error(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (3)
where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{H}}$, $\widehat{\mathbf{T}}$ and $\mathbf{H}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{H}$ and $\mathbf{T}$. In order to prevent the numerical disparity between the two loss terms and stabilize the training for both of the SO(3)-equivariant and SO(3)-invariant branches, we apply $\mu(loss_{H},loss_{T})$, i.e., a coefficient to regularize the relative scale of the two loss terms:
$\mu(loss_{H},loss_{T})=\lambda\cdot No\_Grad(\frac{loss_{H}}{loss_{T}})$ (4)
where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale between the two loss terms and should not itself be a source of training gradients, otherwise it would counteract the gradients from $loss_{T}$ in Eq. (3). All of the encoding and decoding modules are trained jointly by Eq. (3).
@@ 6.1 Experimental Conditions
We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-stable (QS), and QH9-dynamic (QD). The first six databases, consisting of periodic crystalline systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023). The last two databases are from the QH9 benchmark series (Yu et al., 2023a), composed of molecular systems with elements like C, H, O, N and F. These databases present diverse and complex challenges to a regression model. Regarding $MG$, $MM$, and QD, as their samples are prepared from an temperature environment at three-hundred Kelvin, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction. For $BG$, $BB$, $BT$, and $BS$, the twisted structures, with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets. Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling. For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and testing sets, ensuring that the atom number of samples do not overlap across the three subsets. For the $QD$ database, the ’mol’ strategy provided by Yu et al. (2023a) is applied to split the training, validation, and testing sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets. The ’mol’ and ’ood’ strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms and the temporal trajectories, respectively. Detailed statistic information of these databases can be found in the Appendix E.
Implementation details of our method for experiments on these databases are presented in Appendix F.
@@
We use a comprehensive set of metrics to deeply evaluate the accuracy performance of deep learning electronic-structure Hamiltonian prediction models. On the databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023), we follow Yin et al. (2024) to adopt a set of Mean Absolute Error (MAE) metrics between predicted and ground truth Hamiltonians, including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$). These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices. On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper (Yu et al., 2023a), including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians. $\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.